Write the first four terms of the sequence defined by the explicit formula an = n! / n(n + 1).
Answer by greenestamps(13200)



What do you need help with?

Do what the instructions say....

a(1) = 1!/(1*2) = 1/2
a(2) = 2!/(2*3) = ??
a(3) = ??
a(4) = ??